# APCTP-Pre2023-010

### v2

@@ III.1 Sharp turns
$\mathcal{R}^{\prime}$ $=P_{\mathcal{R}}+2\Omega\mathcal{S}\,,$ (53)
$P_{\mathcal{R}}^{\prime}$ $=-(3-\epsilon+\eta)P_{\mathcal{R}}-\kappa^{2}\mathcal{R}\,,$ (54)
$\mathcal{S}^{\prime}$ $=P_{\mathcal{S}}\,,$ (55)
$P_{\mathcal{S}}^{\prime}$ $=-2\Omega P_{\mathcal{R}}-(3-\epsilon+\eta)P_{\mathcal{S}}-\left[\kappa^{2}+m_{\rm eff}+{1\over 2}\eta\left(3-\epsilon+{1\over 2}\eta\right)+{1\over 2}\eta^{\prime}\right]\mathcal{S}\,,$ (56)
@@
$m_{\rm eff}\equiv\mathcal{M}_{nn}+3\Omega^{2}\,.$ (57)
@@
$\mathcal{R}$ $={\mathcal{R}|_{\rm te}\over\kappa_{\rm te}}\left[\kappa\cos\left(\kappa-\kappa_{\rm te}\right)-\sin\left(\kappa-\kappa_{\rm te}\right)\right]+{P_{\mathcal{R}}|_{\rm te}\over\kappa_{\rm te}^{3}}\left[\left(\kappa-\kappa_{\rm te}\right)\cos\left(\kappa-\kappa_{\rm te}\right)-(1+\kappa\kappa_{\rm te})\sin\left(\kappa-\kappa_{\rm te}\right)\right]\,,$ (58)
@@
$\quad+{\pi\kappa^{3/2}\over 2\kappa_{\rm te}^{3/2}}\left\{\left[\mathcal{S}|_{\rm te}\left(\nu-{3\over 2}\right)+P_{\mathcal{S}}|_{\rm te}\kappa_{\rm te}\right]Y_{\nu}(\kappa_{\rm te})-\mathcal{S}|_{\rm te}\kappa_{\rm te}Y_{\nu-1}(\kappa_{\rm te})\right\}J_{\nu}(\kappa)\,,$ (59)
@@
$\nu\equiv\sqrt{{9\over 4}-\mathcal{M}_{\rm nn}}\,,$ (60)
and the subscript “te” denotes the values at the end of the turn. Setting $\kappa=0$ in the above relations gives the asymptotic values for the two perturbations which depend on the the values of the two perturbations and their momenta at the end of the turn. To simplify the analysis we have ignored the slow-roll parameters.
@@
$\Omega={\delta\theta\over\delta N}\left[\Theta\left(N-N_{\rm ti}\right)-\Theta\left(N-N_{\rm te}\right)\right]\,,$ (61)
@@
$X(N_{\rm te})=\left[I+\int^{N_{\rm te}}_{N_{\rm ti}}\,\mathrm{d}N_{1}J(N_{1})+{1\over 2}\int^{N_{\rm te}}_{N_{\rm ti}}\int^{N_{1}}_{N_{\rm ti}}\,\mathrm{d}N_{2}\,\mathrm{d}N_{1}J(N_{1})J(N_{2})+\cdots\right]X_{0}\,.$ (62)
@@
$I+\delta NJ(N_{*})+{1\over 2}(\delta N)^{2}J(N_{*})J(N_{*})+\cdots=\exp\left[J(N_{*})\delta N\right]\,,$ (63)
@@
$\bm{e}_{i}=\begin{pmatrix}1\vspace{0.4em}\\ -\dfrac{\kappa^{2}}{\bar{\lambda}_{i}}\vspace{0.4em}\\ \dfrac{\kappa^{2}+\lambda_{i}\bar{\lambda}_{i}}{2\Omega\bar{\lambda}_{i}}\vspace{0.4em}\\ \lambda_{i}\dfrac{\kappa^{2}+\lambda_{i}\bar{\lambda}_{i}}{2\Omega\bar{\lambda}_{i}}\end{pmatrix}\,,$ (65)
@@
$\lambda_{1,2,3,4}$ $=-{1\over 2}\left[3-\epsilon\pm\sqrt{(3-\epsilon)^{2}-4\kappa^{2}-2m_{\rm eff}\pm 2\sqrt{m_{\rm eff}^{2}+16\kappa^{2}\Omega^{2}}}\right]\,,$ (66)
$\lambda_{1}+\lambda_{2}=\lambda_{3}+\lambda_{4}$ $=-(3-\epsilon)\,,$ (67)
$\bar{\lambda}_{i}$ $\equiv(3-\epsilon)+\lambda_{i}\,.$ (68)
@@
$\mathcal{R}|_{\rm te}$ $=\mathcal{R}^{(\mathcal{R})}|_{\rm te}+\mathcal{R}^{(\mathcal{S})}|_{\rm te}=\mathcal{R}^{(\mathcal{R})}|_{\rm te}\left(\mathcal{R}|_{\rm ti},P_{\mathcal{R}}|_{\rm ti}\right)+\mathcal{R}^{(\mathcal{S})}|_{\rm te}\left(\mathcal{S}|_{\rm ti},P_{\mathcal{S}}|_{\rm ti}\right)\,,$ (69)
$P_{\mathcal{R}}|_{\rm te}$ $=P_{\mathcal{R}}^{(\mathcal{R})}|_{\rm te}+P_{\mathcal{R}}^{(\mathcal{S})}|_{\rm te}=P_{\mathcal{R}}^{(\mathcal{R})}|_{\rm te}\left(\mathcal{R}|_{\rm ti},P_{\mathcal{R}}|_{\rm ti}\right)+P_{\mathcal{R}}^{(\mathcal{S})}|_{\rm te}\left(\mathcal{S}|_{\rm ti},P_{\mathcal{S}}|_{\rm ti}\right)\,,$ (70)
@@
$\quad+{\left(\kappa^{2}-\lambda_{1}\lambda_{2}\right)\left[\lambda_{4}e^{\lambda_{3}\delta N}\left(\lambda_{3}P_{\mathcal{R}}|_{\rm ti}-\kappa^{2}\mathcal{R}|_{\rm ti}\right)-\lambda_{3}e^{\lambda_{4}\delta N}\left(\lambda_{4}P_{\mathcal{R}}|_{\rm ti}-\kappa^{2}\mathcal{R}|_{\rm ti}\right)\right]\over\kappa^{2}(\lambda_{2}-\lambda_{4})(\lambda_{1}-\lambda_{4})(\lambda_{4}-\lambda_{3})}\,,$ (71)
@@
$\quad+{2\Omega\left[\lambda_{4}e^{\lambda_{3}\delta N}\left(P_{\mathcal{S}}|_{\rm ti}-\lambda_{4}\mathcal{S}|_{\rm ti}\right)-\lambda_{3}e^{\lambda_{4}\delta N}\left(P_{\mathcal{S}}|_{\rm ti}-\lambda_{3}\mathcal{S}|_{\rm ti}\right)\right]\over(\lambda_{2}-\lambda_{4})(\lambda_{1}-\lambda_{4})(\lambda_{4}-\lambda_{3})}\,,$ (72)
@@
$\quad+{\left(\kappa^{2}-\lambda_{1}\lambda_{2}\right)\left[e^{\lambda_{3}\delta N}\left(\lambda_{3}P_{\mathcal{R}}|_{\rm ti}-\kappa^{2}\mathcal{R}|_{\rm ti}\right)-e^{\lambda_{4}\delta N}\left(\lambda_{4}P_{\mathcal{R}}|_{\rm ti}-\kappa^{2}\mathcal{R}|_{\rm ti}\right)\right]\over(\lambda_{2}-\lambda_{4})(\lambda_{1}-\lambda_{4})(\lambda_{4}-\lambda_{3})}\,,$ (73)
@@
$\quad+{2\Omega\kappa^{2}\left[e^{\lambda_{3}\delta N}\left(P_{\mathcal{S}}|_{\rm ti}-\lambda_{4}\mathcal{S}|_{\rm ti}\right)-e^{\lambda_{4}\delta N}\left(P_{\mathcal{S}}|_{\rm ti}-\lambda_{3}\mathcal{S}|_{\rm ti}\right)\right]\over(\lambda_{2}-\lambda_{4})(\lambda_{1}-\lambda_{4})(\lambda_{4}-\lambda_{3})}\,.$ (74)
@@
Having found the analytic solutions, now we proceed to compute the power spectrum at the end of inflation. For this goal, we promote the fields to operators and decompose them in terms of the creation and annihilation operators, and solve for the vacuum state mode functions. Then, the power spectrum is obtained by the absolute value squared of the mode function evaluated at the end of inflation. For the current case, the sensible moment to set up the operators is well before the horizon crossing where we can treat both $\mathcal{R}$ and $\mathcal{S}$ quantum mechanically. Thus, the two independent degrees of freedom at early stage are $\mathcal{R}|_{\text{ti}}$ and $\mathcal{S}|_{\text{ti}}$ that can be written in terms of the Bunch-Davies mode function solutions. Consequently, the two asymptotic values for the curvature perturbation (58) for $\kappa\to 0$ can be divided into two parts with one being dependent only on $\mathcal{R}^{(\mathcal{R})}|_{\text{te}}$ and $P_{\mathcal{R}}^{(\mathcal{R})}|_{\text{te}}$, and the other only on $\mathcal{R}^{(\mathcal{S})}|_{\text{te}}$ and $P_{\mathcal{R}}^{(\mathcal{R})}|_{\text{te}}$:
$\mathcal{R}^{(\mathcal{R})}(\kappa\to 0)$ $={\mathcal{R}^{(\mathcal{R})}|_{\rm te}\over\kappa_{\rm te}}\sin\left(\kappa_{\rm te}\right)+{P_{\mathcal{R}}^{(\mathcal{R})}|_{\rm te}\over\kappa_{\rm te}^{3}}\left[\sin\left(\kappa_{\rm te}\right)-\kappa_{\rm te}\cos\left(\kappa_{\rm te}\right)\right]\,,$ (75)
$\mathcal{R}^{(\mathcal{S})}(\kappa\to 0)$ $={\mathcal{R}^{(\mathcal{S})}|_{\rm te}\over\kappa_{\rm te}}\sin\left(\kappa_{\rm te}\right)+{P_{\mathcal{R}}^{(\mathcal{S})}|_{\rm te}\over\kappa_{\rm te}^{3}}\left[\sin\left(\kappa_{\rm te}\right)-\kappa_{\rm te}\cos\left(\kappa_{\rm te}\right)\right]\,,$ (76)
@@
$\mathcal{P}_{\mathcal{R}}={k^{3}\over 2\pi^{2}}\left[\left|\mathcal{R}^{(\mathcal{R})}(\kappa\to 0)\right|^{2}+\left|\mathcal{R}^{(\mathcal{S})}(\kappa\to 0)\right|^{2}\right]\,.$ (77)
@@
Figure 1: The evolution of the curvature perturbation (normalized by the single-field expression $\mathcal{P}_{0}$) for a sharp turn (taking the form of a top-hat function) of total swapped angle $4\pi$ and duration $\delta N=0.1$ $e$-folds. The effective mass has been chosen as $m_{\rm eff}=2\Omega^{2}+5$ and the turn is centred around $N=0$. The two curves correspond to the numerical solution (solid) and the approximate analytic one (dashed) given by the series of formulae presented in section III.1. The negligible difference after the horizon-crossing point ($N\sim 5$) is due to neglecting the slow-roll parameters in the analytic approximation.
III.2 Broad turns
@@
When the turn is sufficiently long in duration, the situation is drastically different. Looking at the formal expression (62), it becomes clear that in principle one has to use the full Dyson series to obtain a reliable solution, [^5] The only exception is the very restricted class of time-dependent matrices $J$ with constant eigenvectors which allows for the simple solution of the form $\exp\left(\int\,\mathrm{d}NJ\right)$. though in most cases some perturbative scheme is assumed and the Dyson series is truncated at a given order. In this subsection we will show how this is done in the case of constant and large turns.
@@
The first (approximate) analytic solution in the case of constant turns was presented in Bjorkmo:2019qno under certain assumptions. First, it relied on the effective field theory description of Garcia-Saenz:2018vqf which predicted a power spectrum at the end of inflation given as $\mathcal{P}=\mathcal{P}_{0}e^{2c}$, with $\mathcal{P}_{0}$ being the single-field power spectrum and $c$ an unknown constant that is derived from the full two-field theory. Second, the Hubble friction terms were assumed negligible on subhorizon scales, an assumption that was also present in predecessor works on the effective field theory for constant turns, e.g. Achucarro:2012yr . Lastly, by pointing out the stability matrix of the coupled $\mathcal{R}$-$\mathcal{S}$ differential system (without the friction terms) acquires a positive eigenvalue at a given time before horizon crossing, the unknown constant $c$ is identified as $\int\,\mathrm{d}N\lambda_{4}$, where $\lambda_{4}$ corresponds to the eigenvalue that becomes positive. Even though the solution was obtained as an approximation it matched numerics to a great accuracy.
@@
However, examining the role of the Hubble friction terms in more detail leads to an apparent contradiction. If the Hubble friction terms are subdominant then on similar grounds one might expect that including those terms into the definition of the fourth eigenvalue would yield more accurate results. Surprisingly though, integrating the more accurate version of the fourth eigenvalue [which we have listed in (66)] severely underestimates the overall amplification by an amount that is beyond the expected accuracy of the method. This happens because the eigenvalue including the friction terms has the form $-(3-\epsilon)+\sqrt{(3-\epsilon)^{2}+A}$, in contrast to the one without the friction terms $\sqrt{A}$ and thus falls off faster as $A$ decreases, i.e. as the mode approaches the horizon, leading to significantly less amplification. This also has the implication that for small $\kappa$ the dominant $\kappa$-dependent part of $\log\mathcal{P}$ takes the form $|c_{\rm s}^{2}|\kappa^{2}$ for the former case, with $c_{\rm s}^{2}$ being the sound speed in the effective description, as opposed to $|c_{\rm s}|\kappa$ for the latter case which is also the correct result. Therefore, to derive the correct formula one has to omit the Hubble friction terms, which seems self-contradicting.
@@
The role of the friction terms is important even for subhorizon scales as they control the amplitude of oscillations. These terms cannot be neglected when solving the equations for an extended period of time. For instance, ignoring Hubble friction in the equation for the single-field curvature perturbation yields the wrong result as it predicts $|\mathcal{R}|\rightarrow 0$ for $N\rightarrow-\infty$ instead of the correct value $|\mathcal{R}|\sim e^{-N}$. Even though a case could be made that these terms would yield subleading corrections for large turns, in reality for this to hold one should consider extreme turns of the order $\mathcal{O}(300)$, well outside the validity of perturbation theory. In what follows we will derive an approximate solution for $\mathcal{R}$ as well as for $\mathcal{S}$, which was not derived in the previous references, without resorting to an effective field theory approach. This will be done by converting the equations for perturbations into perturbative expansion in terms of $\kappa^{-1}$ and find a consistent solution for the zeroth-order part, essentially applying the adiabatic theorem 1950JPSJ….5..435K adjusted to linear differential equations. Moreover, we clarify the role of the Hubble friction terms which, when taken into account, amount for the correct amplitude at horizon crossing.
@@
′′$\mathcal{R}^{\prime\prime}+3\mathcal{R}^{\prime}+x^{2}\mathcal{R}=0\,,$ (78)
where $x$ is defined in (28) and we are interested in solving it in the subhorizon regime, i.e. $x\gg 1$. To this end, we perform the change of variables from $N$ to $x$ to obtain
$\mathcal{R}_{,xx}-{2\over x}\mathcal{R}_{,x}+\mathcal{R}=0\,,$ (79)
@@
$X\equiv\begin{pmatrix}\mathcal{R}\\ p_{\mathcal{R}}\end{pmatrix}\,,\qquad J_{(0)}\equiv\begin{pmatrix}0&1\\ -1&0\end{pmatrix}\,,\qquad J_{(1)}\equiv\begin{pmatrix}0&0\\ 0&\dfrac{2}{x}\end{pmatrix}\,.$ (81)
The momentum with $x$ as the time variable, $p_{\mathcal{R}}$, is related to $P_{\mathcal{R}}$, i.e. the momentum in terms of the $e$-folding time, via the chain rule $P_{\mathcal{R}}=-x\,p_{\mathcal{R}}$. The next step is to diagonalize the zeroth-order stability matrix as $J_{\rm D}\equiv U^{-1}J_{(0)}U$ and rewrite (80) in terms of a new variable $X\equiv UY$:

### v3

@@ 3.1 Sharp turns
$\mathcal{R}^{\prime}$ $=P_{\mathcal{R}}+2\Omega\mathcal{S}\,,$ (3.1)
$P_{\mathcal{R}}^{\prime}$ $=-(3-\epsilon+\eta)P_{\mathcal{R}}-\kappa^{2}\mathcal{R}\,,$ (3.2)
$\mathcal{S}^{\prime}$ $=P_{\mathcal{S}}\,,$ (3.3)
$P_{\mathcal{S}}^{\prime}$ $=-2\Omega P_{\mathcal{R}}-(3-\epsilon+\eta)P_{\mathcal{S}}-\left[\kappa^{2}+m_{\rm eff}+{1\over 2}\eta\left(3-\epsilon+{1\over 2}\eta\right)+{1\over 2}\eta^{\prime}\right]\mathcal{S}\,,$ (3.4)
@@
$m_{\rm eff}\equiv\mathcal{M}_{nn}+3\Omega^{2}\,.$ (3.5)
@@
$\mathcal{R}$ $={\mathcal{R}|_{\rm te}\over\kappa_{\rm te}}\left[\kappa\cos\left(\kappa-\kappa_{\rm te}\right)-\sin\left(\kappa-\kappa_{\rm te}\right)\right]+{P_{\mathcal{R}}|_{\rm te}\over\kappa_{\rm te}^{3}}\left[\left(\kappa-\kappa_{\rm te}\right)\cos\left(\kappa-\kappa_{\rm te}\right)-(1+\kappa\kappa_{\rm te})\sin\left(\kappa-\kappa_{\rm te}\right)\right]\,,$ (3.6)
@@
$\quad+{\pi\kappa^{3/2}\over 2\kappa_{\rm te}^{3/2}}\left\{\left[\mathcal{S}|_{\rm te}\left(\nu-{3\over 2}\right)+P_{\mathcal{S}}|_{\rm te}\kappa_{\rm te}\right]Y_{\nu}(\kappa_{\rm te})-\mathcal{S}|_{\rm te}\kappa_{\rm te}Y_{\nu-1}(\kappa_{\rm te})\right\}J_{\nu}(\kappa)\,,$ (3.7)
@@
$\nu\equiv\sqrt{{9\over 4}-\mathcal{M}_{\rm nn}}\,,$ (3.8)
and the subscript “te” denotes the values at the end of the turn. Setting $\kappa=0$ in the above relations gives the asymptotic values for the two perturbations which depend on the values of the two perturbations and their momenta at the end of the turn. To simplify the analysis we have ignored the slow-roll parameters.
@@
$\Omega={\delta\theta\over\delta N}\left[\Theta\left(N-N_{\rm ti}\right)-\Theta\left(N-N_{\rm te}\right)\right]\,,$ (3.9)
@@
$X(N_{\rm te})=\left[I+\int^{N_{\rm te}}_{N_{\rm ti}}\,\mathrm{d}N_{1}J(N_{1})+{1\over 2}\int^{N_{\rm te}}_{N_{\rm ti}}\int^{N_{\rm te}}_{N_{\rm ti}}\,\mathrm{d}N_{2}\,\mathrm{d}N_{1}J(N_{1})J(N_{2})+\cdots\right]X_{0}\,.$ (3.10)
@@
$I+\delta NJ(N_{*})+{1\over 2}(\delta N)^{2}J(N_{*})J(N_{*})+\cdots=\exp\left[J(N_{*})\delta N\right]\,,$ (3.11)
@@
$\bm{e}_{i}=\begin{pmatrix}1\vspace{0.4em}\\ -\dfrac{\kappa^{2}}{\bar{\lambda}_{i}}\vspace{0.4em}\\ \dfrac{\kappa^{2}+\lambda_{i}\bar{\lambda}_{i}}{2\Omega\bar{\lambda}_{i}}\vspace{0.4em}\\ \lambda_{i}\dfrac{\kappa^{2}+\lambda_{i}\bar{\lambda}_{i}}{2\Omega\bar{\lambda}_{i}}\end{pmatrix}\,,$ (3.13)
@@
$\lambda_{1,2,3,4}$ $=-{1\over 2}\left[3-\epsilon\pm\sqrt{(3-\epsilon)^{2}-4\kappa^{2}-2m_{\rm eff}\pm 2\sqrt{m_{\rm eff}^{2}+16\kappa^{2}\Omega^{2}}}\right]\,,$ (3.14)
$\lambda_{1}+\lambda_{2}=\lambda_{3}+\lambda_{4}$ $=-(3-\epsilon)\,,$ (3.15)
$\bar{\lambda}_{i}$ $\equiv(3-\epsilon)+\lambda_{i}\,.$ (3.16)
@@
$\mathcal{R}|_{\rm te}$ $=\mathcal{R}^{(\mathcal{R})}|_{\rm te}+\mathcal{R}^{(\mathcal{S})}|_{\rm te}=\mathcal{R}^{(\mathcal{R})}|_{\rm te}\left(\mathcal{R}|_{\rm ti},P_{\mathcal{R}}|_{\rm ti}\right)+\mathcal{R}^{(\mathcal{S})}|_{\rm te}\left(\mathcal{S}|_{\rm ti},P_{\mathcal{S}}|_{\rm ti}\right)\,,$ (3.17)
$P_{\mathcal{R}}|_{\rm te}$ $=P_{\mathcal{R}}^{(\mathcal{R})}|_{\rm te}+P_{\mathcal{R}}^{(\mathcal{S})}|_{\rm te}=P_{\mathcal{R}}^{(\mathcal{R})}|_{\rm te}\left(\mathcal{R}|_{\rm ti},P_{\mathcal{R}}|_{\rm ti}\right)+P_{\mathcal{R}}^{(\mathcal{S})}|_{\rm te}\left(\mathcal{S}|_{\rm ti},P_{\mathcal{S}}|_{\rm ti}\right)\,,$ (3.18)
@@
$\quad+{\left(\kappa^{2}-\lambda_{1}\lambda_{2}\right)\left[\lambda_{4}e^{\lambda_{3}\delta N}\left(\lambda_{3}P_{\mathcal{R}}|_{\rm ti}-\kappa^{2}\mathcal{R}|_{\rm ti}\right)-\lambda_{3}e^{\lambda_{4}\delta N}\left(\lambda_{4}P_{\mathcal{R}}|_{\rm ti}-\kappa^{2}\mathcal{R}|_{\rm ti}\right)\right]\over\kappa^{2}(\lambda_{2}-\lambda_{4})(\lambda_{1}-\lambda_{4})(\lambda_{4}-\lambda_{3})}\,,$ (3.19)
@@
$\quad+{2\Omega\left[\lambda_{4}e^{\lambda_{3}\delta N}\left(P_{\mathcal{S}}|_{\rm ti}-\lambda_{4}\mathcal{S}|_{\rm ti}\right)-\lambda_{3}e^{\lambda_{4}\delta N}\left(P_{\mathcal{S}}|_{\rm ti}-\lambda_{3}\mathcal{S}|_{\rm ti}\right)\right]\over(\lambda_{2}-\lambda_{4})(\lambda_{1}-\lambda_{4})(\lambda_{4}-\lambda_{3})}\,,$ (3.20)
@@
$\quad+{\left(\kappa^{2}-\lambda_{1}\lambda_{2}\right)\left[e^{\lambda_{3}\delta N}\left(\lambda_{3}P_{\mathcal{R}}|_{\rm ti}-\kappa^{2}\mathcal{R}|_{\rm ti}\right)-e^{\lambda_{4}\delta N}\left(\lambda_{4}P_{\mathcal{R}}|_{\rm ti}-\kappa^{2}\mathcal{R}|_{\rm ti}\right)\right]\over(\lambda_{2}-\lambda_{4})(\lambda_{1}-\lambda_{4})(\lambda_{4}-\lambda_{3})}\,,$ (3.21)
@@
$\quad+{2\Omega\kappa^{2}\left[e^{\lambda_{3}\delta N}\left(P_{\mathcal{S}}|_{\rm ti}-\lambda_{4}\mathcal{S}|_{\rm ti}\right)-e^{\lambda_{4}\delta N}\left(P_{\mathcal{S}}|_{\rm ti}-\lambda_{3}\mathcal{S}|_{\rm ti}\right)\right]\over(\lambda_{2}-\lambda_{4})(\lambda_{1}-\lambda_{4})(\lambda_{4}-\lambda_{3})}\,.$ (3.22)
@@
Having found the analytic solutions, we now proceed to compute the power spectrum at the end of inflation. For this goal, we promote the fields to operators and decompose them in terms of the creation and annihilation operators, and solve for the vacuum state mode functions. Then, the power spectrum is obtained by the absolute value squared of the mode function evaluated at the end of inflation. For the current case, the sensible moment to set up the operators is well before the horizon crossing where we can treat both $\mathcal{R}$ and $\mathcal{S}$ quantum mechanically. Thus, the two independent degrees of freedom at early stage are $\mathcal{R}|_{\text{ti}}$ and $\mathcal{S}|_{\text{ti}}$ that can be written in terms of the Bunch-Davies mode function solutions. Consequently, the two asymptotic values for the curvature perturbation (3.6) for $\kappa\to 0$ can be divided into two parts with one being dependent only on $\mathcal{R}^{(\mathcal{R})}|_{\text{te}}$ and $P_{\mathcal{R}}^{(\mathcal{R})}|_{\text{te}}$, and the other only on $\mathcal{R}^{(\mathcal{S})}|_{\text{te}}$ and $P_{\mathcal{R}}^{(\mathcal{R})}|_{\text{te}}$:
$\mathcal{R}^{(\mathcal{R})}(\kappa\to 0)$ $={\mathcal{R}^{(\mathcal{R})}|_{\rm te}\over\kappa_{\rm te}}\sin\left(\kappa_{\rm te}\right)+{P_{\mathcal{R}}^{(\mathcal{R})}|_{\rm te}\over\kappa_{\rm te}^{3}}\left[\sin\left(\kappa_{\rm te}\right)-\kappa_{\rm te}\cos\left(\kappa_{\rm te}\right)\right]\,,$ (3.23)
$\mathcal{R}^{(\mathcal{S})}(\kappa\to 0)$ $={\mathcal{R}^{(\mathcal{S})}|_{\rm te}\over\kappa_{\rm te}}\sin\left(\kappa_{\rm te}\right)+{P_{\mathcal{R}}^{(\mathcal{S})}|_{\rm te}\over\kappa_{\rm te}^{3}}\left[\sin\left(\kappa_{\rm te}\right)-\kappa_{\rm te}\cos\left(\kappa_{\rm te}\right)\right]\,,$ (3.24)
@@
$\mathcal{P}_{\mathcal{R}}={k^{3}\over 2\pi^{2}}\left[\left|\mathcal{R}^{(\mathcal{R})}(\kappa\to 0)\right|^{2}+\left|\mathcal{R}^{(\mathcal{S})}(\kappa\to 0)\right|^{2}\right]\,.$ (3.25)
@@
Figure 1: The evolution of the curvature perturbation (normalized by the single-field expression $\mathcal{P}_{0}$) for a sharp turn (taking the form of a top-hat function) of total swapped angle $4\pi$ and duration $\delta N=0.1$ $e$-folds. The effective mass has been chosen as $m_{\rm eff}=2\Omega^{2}+5$ and the turn is centred around $N=0$. The two curves correspond to the numerical solution (solid) and the approximate analytic one (dashed) given by the series of formulae presented in section 3.1. The negligible difference after the horizon-crossing point ($N\sim 5$) is due to neglecting the slow-roll parameters in the analytic approximation.
3.2 Broad turns
@@
When the turn is sufficiently long in duration, the situation is drastically different. Looking at the formal expression (3.10), it becomes clear that in principle one has to use the full Dyson series to obtain a reliable solution, [^5] The only exception is the very restricted class of time-dependent matrices $J$ with constant eigenvectors which allows for the simple solution of the form $\exp\left(\int\,\mathrm{d}NJ\right)$. though in most cases some perturbative scheme is assumed and the Dyson series is truncated at a given order. In this subsection we will show how this is done in the case of constant and large turns.
@@
The first (approximate) analytic solution in the case of constant turns was presented in [59] under certain assumptions. First, it relied on the effective field theory description of [60] which predicted a power spectrum at the end of inflation given as $\mathcal{P}=\mathcal{P}_{0}e^{2c}$, with $\mathcal{P}_{0}$ being the single-field power spectrum and $c$ an unknown constant that is derived from the full two-field theory. Second, the Hubble friction terms were assumed negligible on subhorizon scales, an assumption that was also present in predecessor works on the effective field theory for constant turns, e.g. [61]. Lastly, by pointing out the stability matrix of the coupled $\mathcal{R}$-$\mathcal{S}$ differential system (without the friction terms) acquires a positive eigenvalue at a given time before horizon crossing, the unknown constant $c$ is identified as $\int\,\mathrm{d}N\lambda_{4}$, where $\lambda_{4}$ corresponds to the eigenvalue that becomes positive. Even though the solution was obtained as an approximation it matched numerics to a great accuracy.
@@
However, examining the role of the Hubble friction terms in more detail leads to an apparent contradiction. If the Hubble friction terms are subdominant then on similar grounds one might expect that including those terms into the definition of the fourth eigenvalue would yield more accurate results. Surprisingly though, integrating the more accurate version of the fourth eigenvalue [which we have listed in (3.14)] severely underestimates the overall amplification by an amount that is beyond the expected accuracy of the method. This happens because the eigenvalue including the friction terms has the form $-(3-\epsilon)+\sqrt{(3-\epsilon)^{2}+A}$, in contrast to the one without the friction terms $\sqrt{A}$ and thus falls off faster as $A$ decreases, i.e. as the mode approaches the horizon, leading to significantly less amplification. This also has the implication that for small $\kappa$ the dominant $\kappa$-dependent part of $\log\mathcal{P}$ takes the form $|c_{\rm s}^{2}|\kappa^{2}$ for the former case, with $c_{\rm s}^{2}$ being the sound speed in the effective description, as opposed to $|c_{\rm s}|\kappa$ for the latter case which is also the correct result. Therefore, to derive the correct formula one has to omit the Hubble friction terms, which seems self-contradicting.
@@
The role of the friction terms is important even for subhorizon scales as they control the amplitude of oscillations. These terms cannot be neglected when solving the equations for an extended period of time. For instance, ignoring Hubble friction in the equation for the single-field curvature perturbation yields the wrong result as it predicts $|\mathcal{R}|\rightarrow 0$ for $N\rightarrow-\infty$ instead of the correct value $|\mathcal{R}|\sim e^{-N}$. Even though a case could be made that these terms would yield subleading corrections for large turns, in reality for this to hold one should consider extreme turns of the order $\mathcal{O}(300)$, well outside the validity of perturbation theory. In what follows we will derive an approximate solution for $\mathcal{R}$ as well as for $\mathcal{S}$, which was not derived in the previous references, without resorting to an effective field theory approach. This will be done by converting the equations for perturbations into perturbative expansion in terms of $\kappa^{-1}$ and find a consistent solution for the zeroth-order part, essentially applying the adiabatic theorem [62] adjusted to linear differential equations. Moreover, we clarify the role of the Hubble friction terms which, when taken into account, amount for the correct amplitude at horizon crossing.
@@
′′$\mathcal{R}^{\prime\prime}+3\mathcal{R}^{\prime}+x^{2}\mathcal{R}=0\,,$ (3.26)
where $x$ is defined in (2.28) and we are interested in solving it in the subhorizon regime, i.e. $x\gg 1$. To this end, we perform the change of variables from $N$ to $x$ to obtain
$\mathcal{R}_{,xx}-{2\over x}\mathcal{R}_{,x}+\mathcal{R}=0\,,$ (3.27)
@@
$X\equiv\begin{pmatrix}\mathcal{R}\\ p_{\mathcal{R}}\end{pmatrix}\,,\qquad J_{(0)}\equiv\begin{pmatrix}0&1\\ -1&0\end{pmatrix}\,,\qquad J_{(1)}\equiv\begin{pmatrix}0&0\\ 0&\dfrac{2}{x}\end{pmatrix}\,.$ (3.29)
The momentum with $x$ as the time variable, $p_{\mathcal{R}}$, is related to $P_{\mathcal{R}}$, i.e. the momentum in terms of the $e$-folding time, via the chain rule $P_{\mathcal{R}}=-x\,p_{\mathcal{R}}$. The next step is to diagonalize the zeroth-order stability matrix as $J_{\rm D}\equiv U^{-1}J_{(0)}U$ and rewrite (3.28) in terms of a new variable $X\equiv UY$: